P10-8 Calculating Salvage Value [LO1]

An asset used in a 4-year project falls in the 5-year MACRS class (MACRS Table) for tax purposes. The asset has an acquisition cost of $12,420,000 and will be sold for $2,760,000 at the end of the project.

  

Required:

If the tax rate is 32 percent, what is the aftertax salvage value of the asset?
Answer #1

Year

MACRS %

Depreciation

Accumulated Depreciation

Book Value

1

0.2

2484000

2484000

9936000

2

0.32

3974400

6458400

5961600

3

0.192

2384640

8843040

3576960

4

0.1152

1430784

10273824

2146176

After 4 year book value of the asset is $2146176.

It will be sold for 2760000

Selling price is more than book value so the company will gain on the asset sale

Gain = 2760000 – 2146176 = $613824

Tax will be levied on the asset sale

Tax = 613824* .32 = 196423.68

After tax cash flow from asset sale = 2760000 - 196423.68 = 2563576.3

--------------------------------------------------------------------------------------------------------------------------

Feel free to comment if you need further assistance J

Pls rate this answer if you found it useful.